Project

General

Profile

Actions

HallMeeting-2014-02-19

1. Muon Beamline – Henry Nebrensky

Decay solenoid:

Critical flowmeter is broken, needs to be refurbished or replaced. Can possibly be bridged out to allow cooldown – Mike consult Adrian Oates
Relatively sudden factor ten increase in RGA reading – possibly compressor seal again – Mike – if it is, all bets are off until HPC are consulted.

Conventional magnets:

Q8 PSU handle fitted
Target:
Send irradiated fibres to Sheffield – Andy - stands
Offline target inspection underway today
Online run with frame raised completed: mechanism works fine, control software crashed – Henry consult Matt Robinson

Proton absorber:

Being measured by Metrology, will be issued as MICE note
RAMS for MICE Hall installation started1.
2. FC, liquid hydrogen and absorber integration - Steve Watson

Absorber removed, still need Steve & Andy G to look at broken electrical connector

3. Tracker activity in MICE Hall and R1/R9 – Craig Macwaters

Measuring dummy today, adjustment to mounting cams required, re-install tent base tomorrow - Geoff.
‘Hands-on’ meeting each Wednesday in R9 after this meeting at 14:30 sharp to discuss parallel FC/tracker activity for following week. Minimally, Tracker and FC nominated persons must attend

4. Installation of Step IV – Andy Nichols.
West mezzanine completion – 7st March
Jason – request trench cable repositioning through ISIS Footprints – stands,
also find spec of ISIS control cables and susceptibility to magnet quench – Field Mitigation Group – important action! - stands
Remove tracker readout cryostat from R5.2 ASAP – Mark, when scaffolding dismantled
Brass interlayer pads fitted to downstream SS support base – Andy order greasers
South ramp cutting effectively starts tomorrow

5. Hall Infrastructure – Andy Nichols

RR2 cable channels have been widened
RR2 raised floor – know how to modify cable trays – floor work starts second week March
Transfer responsibility for air supply from Tim Hayler – Andy - Stands
Make neutron monitor signs as per Paul Wright instructions – Andy
Calibrate neutron monitor at Easter shutdown, have spare PSU – Frances Burge
Beamstopper, slipping downwards, MOM to take note and Andy contact mfrs
East end A/C unit discharged, unit hope to be back on by Friday
Need to demonstrate on Hall layout handling of three*500l LHe dewars – Jason/Paul
Crane mods start date confirmed 7th March – cranes will be permitted out for two weeks from that date

6. Alignment and survey –

Tim Hayler Proton absorber survey – Steve, requested

7. Electrical Infrastructure – Andy Gallagher

Green text: action complete, delete from list
Blue text: waiting for information/parts
Red text: normal work in progress
• Install power feeds to RF rack positions. Cables ordered. Drawing issued.
Mark up Rack positions and have floor plates cut then install supply cables. On Hold for STEP IV installation
• Re-position 208v Transformer further east in the trench. Re-position into RR2.
• Re-position D13 on North mezz extn. Waiting on Craig’s rack position changes.
Install network cables & limit cable for Proton Absorber.
Remove webcam from west wall when access is available from scaffold.
R9 Dark Room Lights?
Install cable tray to RR2
Install distribution board in RR2 when access is available.
Install cable tray in RR2 when steel plinths are in position.
EMR rack – install filtered air inlet and reverse fans.
Install telephone and network to R9 Workshop Clean Room.
Add power to webcam on North mezzanine to list
Mobile phone repeater needs fixing
RF network sockets

8. MLCR, controls and networking, Online – Craig Macwaters/Henry Nebrensky

More keys required for MLCR – Andy – stands
Network connections route to RR2 - possible solution – Craig talk to David P

9. PPS – Craig Macwaters

Potterax door bolts broken – John/Craig – Estates helpdesk? - stands

10. RF progress at DL and RAL – Tim Stanley

Nothing to report ,Tim away

11. Detector preparations and running

Nothing to report

12. R9 activity – Steve Watson

FC#1 now condensing He at twice previous rate
PSU tuning next week, up to 5 amps
Looks like FC#2 problem might be understood – movement of cold mass
LHe dewar delivered tomorrow – move big pallet- Mark
Bottle rack agreed with BPG/SHE at North side of R9 – John, order suitable rack/chain

13. Effective dates of next shutdown and job list –All.

4th April – 6th May 2014, (maintenance days: 26th February, 19th March

Synchrotron phone link
Neutron monitor calibration
Proton absorber installation
Drip trays for quads/dipole
Clean up DSA Target chiller – talk to Paul Hodgson
Modify DS transfer line pumping system for ease of access

14. A.O.B

None

15. Date of next meeting February 26th G-06,R66 13.30 local time

Updated by Govans, John over 8 years ago · 1 revisions